OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = -7 and a6 = -7,168?

OpenStudy (anonymous):

a6 = 168 ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Sorry I can't understand the given ?!

OpenStudy (anonymous):

what given?

OpenStudy (anonymous):

a1 = -7 what else is given to solve this problem ? value of a6 ?

OpenStudy (anonymous):

a6= -7168

OpenStudy (anonymous):

so the GS (a, ar , ar^2 ,...) so a = a1 = -7 a6 = \[ar ^{5}\] -7168=-7 * r ^5 r=4 a9 = ar^8 = (-7)(4^8)= -458752

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

urw :)